OpenStudy (idkwut):

What is the electric force between a +2 µC point charge and a –2 µC point charge if they are separated by a distance of 5.0 cm? Show your work. (µC = 1.0 × 10–6 C)

OpenStudy (praxer):

Do you the know the formula of the electrostatic force between two charges ? @IDKwut

OpenStudy (anonymous):

The equation @praxer is referring to is known as Coulomb's Law.

OpenStudy (idkwut):

I am kinda sorta. Lol, just came back from spring break so please refresh my memory! :) Isn't it F = q1q2/r^2 or something.

OpenStudy (nali):

\[F= Q1 Q2 / 4\pi \epsilon0 r ^{2} \]

OpenStudy (anonymous):

Close! You're forgetting Coulomb's Constant, \(k_e\).

OpenStudy (nali):

actually this is right or you could also use F= k Q1Q2 / r^2 any way is right

OpenStudy (idkwut):

Oh yea. F = kq1q1/r^2 which is equivalent to the formula Nali posted.

OpenStudy (anonymous):

\(k_e\) is equal to \( 1 \over 4 \pi \epsilon_0\)

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

We know both charges and the distance between them. You just need to plug the values in!

OpenStudy (idkwut):

Sorry, I can't really see what you posted (the k = to 1/...) It's rather small.

OpenStudy (praxer):

$$\huge k=\frac{1}{4\pi\epsilon_o}$$

OpenStudy (nali):